OpenStudy (anonymous):

c'=(e^-4t)(1-4t) how do I find the maximum value of "c" and when it occurs "t"

OpenStudy (turingtest):

the maximum occurs when c'=0

OpenStudy (turingtest):

max or min*

OpenStudy (turingtest):

when is e^-4t=0 ?

OpenStudy (anonymous):

when t=0?

OpenStudy (turingtest):

is it hm... as I recall e^0=1, so that's not it...

OpenStudy (anonymous):

oh right. dang.

OpenStudy (turingtest):

what would happen if we tried to take the log of both sides?

OpenStudy (anonymous):

ummm im not sure

OpenStudy (turingtest):

remember that\[e^{-4t}=\frac1{e^{4t}}\]notice that as \(t\to\infty\) we have \(e^{-4t}\to0\), so t would have to be infinity for e^-4t to be zero, so it is \(never\) zero

OpenStudy (turingtest):

that means we only have to solve the possibility that the other part is zero, so solve 1-4t=0

OpenStudy (anonymous):

so t=1/4? and then i plug that back into the original equation?

OpenStudy (anonymous):

well i did have an original equation of c=te^-4t and im supposed to find the max value of "c" and when it occurs "t"

OpenStudy (turingtest):

okay, so the maximum value will be what you get by plugging t into that

OpenStudy (turingtest):

t=1/4 is where the maximum occures, and c(1/4) is the max value

OpenStudy (anonymous):

so to find it i put 1/4 in for t? c=1/4 * e^-4(1/4)

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

ok that makes sense. Thank you so much!

OpenStudy (turingtest):

welcome!